OpenStudy (anonymous):

Please help with step-by-step notes for the laplace transform: y'' +6y'+5y = 12e^t

OpenStudy (anonymous):

initial conditions??

OpenStudy (anonymous):

sorry, y(0) = -1 y'(0) = 7

OpenStudy (zzr0ck3r):

do you know the laplace of dy/dt?

OpenStudy (anonymous):

F(s)?

OpenStudy (zzr0ck3r):

L(dy/dt) = sL(y)-y(0)

OpenStudy (anonymous):

oh, that formula. yes. i usually can get to the partial fractions part, but I usually have trouble separating the terms and whatnot

OpenStudy (zzr0ck3r):

L(d^2y/dt^2) = s^2L(y)-y`(0) - sy(0)

OpenStudy (zzr0ck3r):

well show me what you got when you get to the partial fractions part

OpenStudy (anonymous):

Y(s) = (12 -s+13)/((s-1)(s^2 +6s+5))

OpenStudy (anonymous):

and that is where i am confused

OpenStudy (anonymous):

....and that's wrong...

OpenStudy (anonymous):

now i have Y(s) = (-s^2 + 14s -13) / ((s-1)(s^2 +6s +5))

OpenStudy (zzr0ck3r):

1 sec

OpenStudy (anonymous):

thank you

OpenStudy (zzr0ck3r):

let me know if you need more help

OpenStudy (anonymous):

thank you

OpenStudy (zzr0ck3r):

np

OpenStudy (anonymous):

oh snap. matrices?

OpenStudy (zzr0ck3r):

yeah do u know them?

OpenStudy (zzr0ck3r):

you can do the same thing just solving three equations with three unknowns, its just longer.

OpenStudy (zzr0ck3r):

are you getting it?